Technical Staff Blog

Last update on .

The filesystem and all services have been restored and are back to normal.  Our efforts to repair file corruption in the file system were not entirely successful.  Following instructions from IBM, we made four attempts to execute their file repair program, and it failed or hung each time.  However, each run did make progress, and ...

Last update on .

A virtual machine server crashed this morning at around 5am, affecting 49 CS Department servers, including list, web, and other important services.  The technical staff spent much of today locating and fixing problems that arose from this event.  All or most services have been restored, but problem mail was interrupted mid-day today, and remains broken ...

Last update on .

During the final weeks of the semester, the technical staff encountered issues with routine maintenance of our GPFS filesystem.  IBM suggested that there was corruption in the filesystem, and recommended that we take it offline to repair it.  We postponed that work until after commencement, to avoid disrupting the final weeks of classes, a conference ...

Last update on .

(copied from John Bazik's email 11:03pm Thursday, May 30th)

The CS filesystem will be shut down for a period of time this weekend. 
This will render desktop and grid systems unusable and disable some
services.  The technical staff is working with IBM to resolve a serious
issue, and this work cannot wait.

The ...

Last update on .

Shortly after 10am this morning, most CS services, including the filesystem, webserver, and list server were affected by the failure of several of the University's virtual machine hosts.  Services were down for about a half hour.  At this time, the cause of the failure is unknown, but we are investigating.  All services were restored ...

Last update on .

Emergency Linux Workstation Reboot

Tstaff will be rebooting several department-managed Linux workstations tomorrow, Wednesday morning at 5am. This is part of an ongoing effort to track down some recent filesystem-related problems.

Workstations to be rebooted include all machines in the Sunlab and some other undergrad spaces, in addition to a few faculty and grad student machines. The complete list ...

Last update on .

CIS has reported that this issue has been resolved. If you are still experiencing issues please contact them at the number below.

 

CIS has reported that a significant amount of users are experiencing an outage of Brown email (Gmail). More info from CIS below:

We are aware of a problem with Google Email - Brown Gmail ...

Last update on .

The department list server stopped delivering mail starting at least as far back as yesterday morning (Thursday the 10th). The problem has been resolved and the list server is chewing through the backlog of messages. If you continue to have problems sending email to department lists or receiving listserv messages, email problem.

Yearly archives